OpenStudy (anonymous):

et of all triples of real numbers with standard vector addition but with scalar multiplication defined by k(x,y,z)=(k^2,k^y,k^z) axiom 8 doesn't hold but I don't see how (axiom 8 is (k+m)u=ku+km) which seems like it holds to me

OpenStudy (p0sitr0n):

it is distributivity. Take u in R^3, m,k in R (k+m)u=((k+m)^2,(k+m)^y,(k+m)^z) although to prove it is a ring (or vector space) you would have to use the binomial expansion, it is only enough to give a counter example to disprove the statement. we see that the first coordinate must always be something^2 (k+m)^2=k^2+2km+m^2 but we see that we can arbitrary choose k and m so that they are not powers of two (like take 3 and 5). Thus the distributivity fails.

OpenStudy (p0sitr0n):

sorry, squares, not powers of 2

OpenStudy (anonymous):

well I see it like this: if we solve the LHS: (k+m)^2u1, (k+m)^2u2,(k+m)u^3 RHS: k^2u1, k^2u2,k^2u3 + m^2u1,m^2u2,m^2u3 --> k^2u1+m^2u1,m^2u2+k^2u2,k^2u3+m^2u3 --> u1(k^2+m^2) ANDDDDD... isee where I went wrong... you cant do (k^2+m^2)=(k+m)^2. Thank you kind stranger!